# Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control

### v4

@@ 1 Introduction
In recent years, deep reinforcement learning (DRL) has shown remarkable success in finding optimal policies across various domains, including games [1, 2] and control tasks [3, 4]. In particular, various model-free DRL algorithms, such as, deep deterministic policy gradient (DDPG) [3], twin delayed DDPG (TD3) [5], soft actor-critic (SAC) [6], and proximal policy optimization (PPO) [7], have demonstrated the ability to generate effective control inputs without requiring explicit system models. However, a critical challenge arises from the discrepancy between the simulation environments used for training and the real-world environment. Therefore, developing DRL policies that are robust to unseen disturbances is crucial for ensuring reliability in real-world applications. To address this challenge, in this paper, we introduce adversarial learning, where the agent is trained in the presence of disturbances. By exposing the policy to worst-case scenarios during training, the agent can improve its robustness and generalization to unseen environments.
In this paper, we present the so-called robust deterministic policy gradeint (RDPG) algorithm which enhances the robustness of DPG algorithm from the concept of the $H_{\infty}$ control and the two-player zero-sum dynamic game perspectives. The $H_{\infty}$ control method is a robust control technique that aims to design an optimal controller that minimizes the impact of disturbances on system performance. A well-established result in control theory shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game, or equivalently, a min-max optimization problem. In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances [8].
Building upon this theoretical foundation, we transform the $H_{\infty}$ control problem into a two-player zero-sum dynamic game between a user agent (controller) and an adversarial agent (disturbance generator). We then propose a reinforcement learning approach based on the deterministic policy gradient (DPG) method to simultaneously train both agents. Each agent learns its own policy: the user aims to minimize the objective function, while the adversary aims to maximize it. As a result of this zero-sum game framework, the user learns a control policy that is inherently robust against worst-case disturbances. Furthermore, we extend this idea to high-dimensional continuous control settings by incorporating it into a deep reinforcement learning framework. The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3.
@@
    We propose the robust deterministic policy gradient (RDPG) algorithm, which enhances robustness against external disturbances by integrating concepts from $H_{\infty}$ control and two-player zero-sum dynamic games into the actor-critic framework.
@@ 2 Related works
The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [9, 10, 11]. However, these methods require precise model linearization and incur high computational costs due to the need to solve Hamilton-Jacobi-Isaacs (HJI) equations. To address these limitations, alternative approaches leverage deep reinforcement learning (DRL) to determine optimal control policies without the need for solving nonlinear HJI equations.
Recent efforts have extended the ideas of $H_{\infty}$ control and two-player zero-sum games into the DRL framework to enhance robustness under various scenarios. For instance, [12] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks. Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework [13], where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward. In contrast, our method explicitly introduces external disturbances into the environment dynamics and formulates the learning process as a two-player zero-sum game between the agent and the adversary. By doing so, our approach enables the agent to learn a control policy that is robust to worst-case disturbance scenarios, rather than merely compensating for action-level noise. In [14], the authors proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator. The disturbance is conditioned on the robot’s state and produces appropriate external forces to facilitate policy learning. To address modeling errors and discrepancies between training and testing conditions, robust adversarial reinforcement learning (RARL) [15] was proposed. RARL trains the agent to perform effectively in the presence of a destabilizing adversary that applies disturbance forces to the system. The agent and the adversary are concurrently trained using the trust region policy optimization (TRPO) algorithm, with the agent learning to maximize the expected reward while the adversary learns to minimize it. Although the underlying concept is similar, their approach differs significantly from ours. Both methods adopt stochastic policy frameworks, such as PPO and TRPO, and they do not impose explicit constraints on the adversary, which may lead to the generation of excessively strong or unrealistic disturbances during training. In contrast, our method is based on deterministic policy gradients and explicitly incorporates the $H_{\infty}$ control perspective into a two-player zero-sum game formulation, enabling more stable and practical learning in continuous control tasks.
@@ 3.1 $H_{\infty}$ Control
where $x_{k}\in\mathbb{R}^{n}$ is the state, $u_{k}\in\mathbb{R}^{m}$ is the control input, $w_{k}\in\mathbb{R}^{p}$ is the disturbance, and $v_{k}\in\mathbb{R}^{q}$ is the process noise at time step $k$. Using the state-feedback controller $u=\pi(x)$, the system can be reduced to the autonomous closed-loop system
@@ 3.2 Two-player zero-sum dynamic game perspective
We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ for the user and $\mu:{\mathbb{R}}^{n}\to{\mathbb{R}}^{p}$ for the adversary
$u=\pi(x)\in{\mathbb{R}}^{m},\quad w=\mu(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}}^{n}$
@@ 4 Proposed method
Figure  1 illustrates the architecture of the proposed method, RDDPG, based on a two-player zero-sum dynamic game. In this method, two policy, user policy $\pi$ and adversary policy $\mu$, interact within the environment. The user policy generates the action $u_{t}$, while the adversary policy generates the disturbance $w_{t}$, both based on the current state $x_{t}$. These inputs are fed into the environment, which returns the next state $x_{t+1}$ and the cost $c_{t}$ based on the dynamics of the environment. The critic is updated based on $\{x_{t},u_{t},w_{t},x_{t+1},c_{t}\}$ and evaluates the joint performance of the user and adversary. The policy of user and adversary are updated by deterministic policy gradient method to minimize and maximize the objective function $J^{\pi,\mu}$. The detailed procedure is described in the following sections.
@@ 4.1 Robust deterministic policy gradient
To apply DPG to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}^{n}}$
where $\pi_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denotes the user’s control policy parameterized by $\theta$ and $\mu_{\phi}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{p}$ denotes the adversary’s control policy parameterized by $\phi$.. The saddle-point problem is then converted to
@@
where $\tilde{c}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{p}\times{\mathbb{R}}^{n}\to{\mathbb{R}}$ is the cost function for the user. Note that the cost function $\tilde{c}$ can be set arbitrarily and is more general than the quadratic cost in the $H_{\infty}$ control problem. This is because the $H_{\infty}$ control problem can be recovered by setting
@@ 4.2 Robust deep deterministic policy gradient
To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3)[5], resulting in the robust deep deterministic policy gradient (RDDPG). In particular, we introduce the following networks:
@@ 4.2.1 Critic update
where $\alpha_{\rm critic}$ is the step-size.
@@ 4.2.2 Actor update
Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite roles.
4.2.3 Target parameters update
The target parameters are updated through the averaging
$\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime},\quad\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
for the actor target parameters and
${\psi_{i}}^{\prime}\leftarrow\tau{\psi_{i}}+(1-\tau){\psi_{i}}^{\prime},\quad i\in\{1,2\}$
for the critic target parameters.
@@ 4.2.4 Exploration
For exploration, we can use
@@
where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises [5]. The overall algorithm is summarized in Algorithm 1.
@@
12:   Store the transition tuple $(x_{k},u_{k},w_{k},c_{k},x_{k+1})$ in the replay buffer $D$
@@
$\theta_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}$
$\phi_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\phi_{i}$
@@ 5.2 State space
where $\mathbf{p}_{k}^{\text{target}}$ and $\mathbf{p}_{k}$ denote the position vectors of the target and the quadrotor, respectively, and $\mathbf{v}_{k}^{\text{target}}$ and $\mathbf{v}_{k}$ represent their corresponding linear velocity vectors. $\phi_{k}$$\theta_{k}$$\psi_{k}$ denotes roll, pitch, yaw of the quadrotor. As in the method used in DQN[2], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
@@ 6.2 Results
Figure 4 illustrates the trajectory tracking performance of each algorithm across varying wind speeds. RDDPG achieves the most accurate trajectory tracking, exhibiting the smallest tracking error compared to the other algorithms. The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor is presented in Figure  5. Across all scenarios, RDDPG consistently achieves lower position errors compared to TD3 and SAC. In the case of DDPG, it exhibits lower position errors under mild wind disturbance scenarios. However, as the disturbance becomes stronger, its performance degrades significantly, resulting in higher errors compared to RDDPG. Furthermore, DDPG shows larger standard deviations, indicating less stable performance across episodes. Detailed numerical results are provided in at Appendix Table  5.
@@ 7 Conclusion
This paper proposes robust deterministic policy gradient (RDPG), which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms. In RDPG, the user seeks to minimize the cost function while the adversary attempts to maximize it, thereby enabling the agent to learn a control policy that is robust to external disturbances. Furthermore, we introduce robust deep deterministic policy gradient (RDDPG), which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks. In order to evaluate the robustness of RDDPG, we implement it on quadrotor tracking tasks under various external disturbances. Experimental results show that RDDPG successfully learns an optimal control policy while maintaining stable performance across different disturbance scenarios. These findings demonstrate that RDDPG outperforms other DRL-based methods in terms of robustness and reliability in real-world-inspired environments.

### v5

# Robust Deterministic Policy Gradient in Zero-Sum Games for Disturbance Attenuation

@@ 1 Introduction
In recent years, deep reinforcement learning (DRL) has shown remarkable success in finding optimal policies across various domains, including games [1, 2] and control tasks [3, 4]. In particular, various model-free DRL algorithms, such as, deep deterministic policy gradient (DDPG) [3], twin delayed DDPG (TD3) [5], soft actor-critic (SAC) [6], and proximal policy optimization (PPO) [7], have demonstrated the ability to generate effective control inputs without requiring explicit system models. However, a critical challenge arises from the discrepancy between the simulation environments used for training and the real-world environment [8, 9, 10]. Therefore, developing DRL policies that are robust to unseen disturbances is crucial for ensuring reliability in real-world applications. To address this challenge, in this paper, we introduce adversarial learning, where the agent is trained in the presence of disturbances. By exposing the policy to worst-case scenarios during training, the agent can improve its robustness and generalization to unseen environments.
In this paper, we present the so-called robust deterministic policy gradeint (RDPG) algorithm which enhances the robustness of deterministic policy gradient (DPG) algorithm [11] from the concept of the $H_{\infty}$ control and the two-player zero-sum dynamic game perspectives. The $H_{\infty}$ control [12, 13, 14] method is a robust control technique that aims to design an optimal controller that minimizes the impact of disturbances on system performance. A well-established result in control theory shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game, or equivalently, a min-max optimization problem [15]. In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances [15].
Building upon this theoretical foundation, we transform the $H_{\infty}$ control problem into a two-player zero-sum dynamic game between the user (the controller) and the adversarial agent. We then propose a reinforcement learning approach based on the DPG method to simultaneously train both agents. Each agent learns its own policy: the user aims to minimize the objective function, while the adversary aims to maximize it. As a result of this zero-sum game framework, the user learns a control policy that is inherently robust against worst-case disturbances. Furthermore, we extend this idea to high-dimensional continuous control settings by incorporating it into a deep reinforcement learning framework. The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3 [5].
@@
    We propose the RDPG algorithm, which enhances robustness against external disturbances by integrating concepts from $H_{\infty}$ control and two-player zero-sum dynamic games into the actor-critic framework.
@@ 2 Related works
The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [14, 16, 17, 18]. However, these methods require precise model linearization and incur high computational costs due to the need to solve Hamilton-Jacobi-Isaacs (HJI) equations. To address these limitations, alternative approaches leverage deep reinforcement learning (DRL) to determine optimal control policies without the need for solving nonlinear HJI equations.
Recent efforts have extended the ideas of $H_{\infty}$ control and two-player zero-sum games into the DRL framework to enhance robustness under various scenarios. For instance, Deshpande et al. [19] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks. Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework [9], where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward. In contrast, our method explicitly introduces external disturbances into the environment dynamics and formulates the learning process as a two-player zero-sum game between the agent and the adversary. By doing so, our approach enables the agent to learn a control policy that is robust to worst-case disturbance scenarios, rather than merely compensating for action-level noise.  Long et al. [10] proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator. The disturbance is conditioned on the robot’s state and produces appropriate external forces to facilitate policy learning. To address modeling errors and discrepancies between training and testing conditions, robust adversarial reinforcement learning (RARL) [8] was proposed. RARL trains the agent to perform effectively in the presence of a destabilizing adversary that applies disturbance forces to the system. The agent and the adversary are concurrently trained using the trust region policy optimization (TRPO) algorithm, with the agent learning to maximize the expected reward while the adversary learns to minimize it. Although the underlying concept is similar, their approach differs significantly from ours. Both methods adopt stochastic policy frameworks, such as PPO and TRPO, and they do not impose explicit constraints on the adversary, which may lead to the generation of excessively strong or unrealistic disturbances during training. In contrast, our method is based on deterministic policy gradients and explicitly incorporates the $H_{\infty}$ control perspective into a two-player zero-sum game formulation, enabling more stable and practical learning in continuous control tasks.
@@ 3.1 $H_{\infty}$ Control
where $x_{k}\in\mathbb{R}^{p}$ is the state, $u_{k}\in\mathbb{R}^{m}$ is the control input, $w_{k}\in\mathbb{R}^{n}$ is the disturbance, and $v_{k}\in\mathbb{R}^{l}$ is the process noise at time step $k$. Using the state-feedback controller $u=\pi(x)$, the system can be reduced to the autonomous closed-loop system
@@ 3.2 Two-player zero-sum dynamic game perspective
We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{p}\to{\mathbb{R}}^{m}$ for the user and $\mu:{\mathbb{R}}^{p}\to{\mathbb{R}}^{n}$ for the adversary
$u=\pi(x)\in{\mathbb{R}}^{m},\quad w=\mu(x)\in{\mathbb{R}}^{n},\quad x\in{\mathbb{R}}^{p}$
@@ 4 Proposed method
Figure 1 illustrates the architecture of the proposed method, RDDPG, based on a two-player zero-sum dynamic game. In this method, two policy, user policy $\pi$ and adversary policy $\mu$, interact within the environment. The user policy generates the action $u_{t}$, while the adversary policy generates the disturbance $w_{t}$, both based on the current state $x_{t}$. These inputs are fed into the environment, which returns the next state $x_{t+1}$ and the cost $c_{t}$ based on the dynamics of the environment. The critic is updated based on $\{x_{t},u_{t},w_{t},x_{t+1},c_{t}\}$ and evaluates the joint performance of the user and adversary. The policy of user and adversary are updated by deterministic policy gradient method to minimize and maximize the objective function $J^{\pi,\mu}$. The detailed procedure is described in the following sections.
@@ 4.1 Robust deterministic policy gradient
To apply deterministic policy gradient (DPG) [11] to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{n},\quad x\in{\mathbb{R}^{p}}$
where $\pi_{\theta}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{m}$ denotes the user’s control policy parameterized by $\theta$ and $\mu_{\phi}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{n}$ denotes the adversary’s control policy parameterized by $\phi$. The saddle-point problem is then converted to
@@
where $\tilde{c}:{\mathbb{R}}^{p}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\times{\mathbb{R}}^{p}\to{\mathbb{R}}$ is the cost function for the user. Note that the cost function $\tilde{c}$ can be set arbitrarily and is more general than the quadratic cost in the $H_{\infty}$ control problem. This is because the $H_{\infty}$ control problem can be recovered by setting
@@ 4.2 Robust deep deterministic policy gradient
To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3)[5], resulting in the RDDPG. In particular, we introduce the following networks:
@@ 4.2.1 Critic update
where $\alpha_{\rm critic}$ is the step-size. After the online critics are updated, the target parameters of two critic $\psi^{\prime}_{1}$ and $\psi^{\prime}_{2}$ are updated as follows:
${\psi^{\prime}_{i}}\leftarrow\tau{\psi_{i}}+(1-\tau){\psi^{\prime}_{i}},\quad i\in\{1,2\}$
where $\tau\in(0,1)$ serves as the interpolation coefficient, facilitating a gradual update that enhances training stability.
@@ 4.2.2 Actor update
Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite roles. After the online actors are updated, the target parameters of actor $\theta^{\prime}$ and $\phi^{\prime}$ are updated as follows
$\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime},\quad\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime}$
where $\tau\in(0,1)$ is the interpolation coefficient.
@@ 4.2.3 Exploration
For exploration during training, Gaussian noise is added to both the user’s control action and the adversary’s disturbance. Specifically, the user’s action $u_{k}$ and the adversary’s disturbance $w_{k}$ at time step $k$ are expressed as
@@
where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ represent zero-mean Gaussian exploration noises with standard deviation $\sigma$ [5]. The overall algorithm is summarized in Algorithm 1.
@@
12:   Store the transition tuple $(x_{k},u_{k},w_{k},c_{k+1},x_{k+1})$ in the replay buffer $D$
@@
$\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
$\phi_{i}^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime}$
@@ 5.2 State space
where $\mathbf{p}_{k}^{\text{target}}$ and $\mathbf{p}_{k}$ denote the position vectors of the target and the quadrotor, respectively, and $\mathbf{v}_{k}^{\text{target}}$ and $\mathbf{v}_{k}$ represent their corresponding linear velocity vectors. $\phi_{k}$$\theta_{k}$$\psi_{k}$ denotes roll, pitch, yaw of the quadrotor. As in the method used in deep Q-network (DQN) [2], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
@@ 6.2 Results
Figure 4 illustrates the trajectory tracking performance of each algorithm across varying wind speeds. RDDPG achieves the most accurate trajectory tracking, exhibiting the smallest tracking error compared to the other algorithms. The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor is presented in Figure 5. Across all scenarios, RDDPG consistently achieves lower position errors compared to TD3 and SAC. In the case of DDPG, it exhibits lower position errors under mild wind disturbance scenarios. However, as the disturbance becomes stronger, its performance degrades significantly, resulting in higher errors compared to RDDPG. Furthermore, DDPG shows larger standard deviations, indicating less stable performance across episodes. Detailed numerical results are provided in at Appendix Table 5.
@@ 7 Conclusion
This paper proposes RDPG, which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms. In RDPG, the user seeks to minimize the cost function while the adversary attempts to maximize it, thereby enabling the agent to learn a control policy that is robust to external disturbances. Furthermore, we introduce RDDOG, which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks. In order to evaluate the robustness of RDDPG, we implement it on quadrotor tracking tasks under various external disturbances. Experimental results show that RDDPG successfully learns an optimal control policy while maintaining stable performance across different disturbance scenarios. These findings demonstrate that RDDPG outperforms other DRL-based methods in terms of robustness and reliability in real-world-inspired environments.